# Scattering polarimetry for correlative nerve fiber imaging and multi-modal analysis

### v2

@@ Acknowledgements
Supplementary Figure 1: Comparison of fiber directions from all modalities. Fiber direction obtained from Mueller polarimetry, 3D-PLI and ComSLI (all three directions displayed in a fiber direction map). The fourth column shows the absolute difference in fiber direction between 3D-PLI and the closest direction obtained by ComSLI $|\phi_{\textrm{PLI}}-\phi_{\textrm{SLI}}|$. Background areas are displayed in black. a Human brain section (high scattering, low transmittance). b Vervet brain section (low scattering, high transmittance).
@@
Supplementary Figure 2: Fourier fit and residuals for exemplary pixels. 3D-PLI measurements with a left-handed circular (LC) and right-handed circular (RC) setting of the PSA are compared. a Fourier fit for the human brain section (low transmittance, high scattering). The two exemplary locations are marked in the transmittance map. b Corresponding residuals and $R^{2}$. c Fourier fit for the vervet monkey brain section (high transmittance, low scattering). The two exemplary locations are marked in the transmittance map. d Corresponding residuals and $R^{2}$ values.

### v3

@@ Acknowledgements
Supplementary Figure 1: Comparison of fiber directions from all modalities: Mueller polarimetry, 3D-PLI and ComSLI (all three directions displayed in a fiber direction map). The fourth column shows the absolute difference in fiber direction between 3D-PLI and the closest direction obtained by ComSLI $|\phi_{\textrm{PLI}}-\phi_{\textrm{SLI}}|$. Background areas are displayed in black. (a) Human brain section (high scattering, low transmittance). (b) Vervet brain section (low scattering, high transmittance).
@@
Supplementary Figure 2: Fourier fit and residuals for exemplary pixels. 3D-PLI measurements with a left-handed circular (LC) and right-handed circular (RC) setting of the PSA are compared. (a) Fourier fit for the human brain section (low transmittance, high scattering). The two exemplary locations are marked in the transmittance map. (b) Corresponding residuals and $R^{2}$. (c) Fourier fit for the vervet monkey brain section (high transmittance, low scattering). The two exemplary locations are marked in the transmittance map. (d) Corresponding residuals and $R^{2}$ values.